How many electrons must be removed from a neutral metal sphere so that it acquire charge +1 nC? If the radius of the sphere is 0.05 m, calculate (i) surface density of charge on the sphere (ii) electric intensity at a point outside and close to the surface of the sphere and (iii) electric intensity at a point of distance 0.08 m from the centre of the sphere

Grade:12

1 Answers

Arun
4 years ago
1 e has charge of 1.6*10^-19
so to get 1nC charge we need to remove =1X10^-9/( 1.6*10^-19) =6.25X10^9 electrons.
Charge density =q/A =1X10^-9/(4*3.14*.05*.05) =31.85nC/m2
 
Hope it helps
 
Regards
Arun